You have been offered the opportunity to invest in a project that will pay $4,856 per year at the end of years one through three and $11,840 per year at the end of years four and five. If the appropriate discount rate is 14.4 percent per year, what is the present value of this cash flow pattern?

Round the answer to two decimal places.
Answer #1

This question requires application of basic time value of money function, according to which EV PV (1 where FV is Future Valu

PV = \frac{FV}{(1 + r)^n}

PV = \frac{4856}{(1 + 0.144)^1} + \frac{4856}{(1 + 0.144)^2} + \frac{4856}{(1 + 0.144)^3} + \frac{11840}{(1 + 0.144)^4} + \frac{11840}{(1 + 0.144)^5}

PV = $4,244.76 + $3,710.45 + $3,243.40 + $6,912.70 + $6,042.57

PV = $24,153.87